OpenStudy (anonymous):

#2 :3 only 6 left to go before my parents say I can go to the concert c: Write the equation of a line in point slope form that has a slope of 4 3 − and passes through the point (4, 5).

OpenStudy (antonio_xx2):

tbh wuts the question

OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5)

OpenStudy (anonymous):

idk why that didnt come through the first time :c lol

OpenStudy (antonio_xx2):

So it would be y-(-10)=3(x-2.5). y+10=3x-7.5 y=3x-17.5 c=-17.5 and i think thats it